# Towards LLM-RecSys Alignment with Textual ID Learning

### v1

@@ Abstract.
Generative recommendation based on Large Language Models (LLMs) have transformed the traditional ranking-based recommendation style into a text-to-text generation paradigm. This approach has attracted significant attention. However, in contrast to standard natural language processing (NLP) tasks that inherently operate on human vocabulary, current research in generative recommendations struggles to effectively encode recommendation items within the text-to-text framework using concise yet meaningful ID representations. Due to this unresolved issue, the potential of LLM-based generative recommendation systems remains largely unexplored.
@@
Recommender System, Information Retrieval, Neural Language Processing
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 14–18, 2024; Washington D.C., USA.price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Natural language generationccs: Computing methodologies Machine learning
@@ 1. Introduction
A few attempts have been made to tackle this problem. P5 (Geng et al., 2022), one of the first works in generative recommendation, proposes allocating out-of-vocabulary (OOV) tokens to items within the recommendation platform. These assigned IDs are fixed-length numerical tokens roughly created based on their sequential appearance in the dataset (e.g., 1001 for the first item, 1002 for the second item). Later, some follow-up research (Hua et al., 2023) evaluates how recommendation performance can be further improved by using different strategies to initialize the numerical IDs.
@@
We note that some recent LLM-based recommendation models employ an encoder-only structure, similar to BERT, as their representation function. Some of these models (Hou et al., 2022; Li et al., 2023a) also possess (or partially possess) characteristics of a foundational recommendation model, though they do not function as generative models. The distinctions among these models are depicted in Table 1. However, generative models present several advantages over discriminative methods. These include transforming the retrieval and ranking processes into a more streamlined generative process, eliminating the need of one-by-one item score calculation, and leveraging the extensive knowledge embedded in pre-trained generative LLMs. Nonetheless, these encoder-only methods remain valuable as baselines for zero-shot evaluation.
@@ 2.3. Diverse ID Generation
The ID generator employs DBS in generating IDs. To ensure the uniqueness of the generated IDs, the algorithm generates $k$ groups of IDs each time and compares each generated ID against a set of already existing IDs. If a duplicate is detected, the algorithm increases the diversity penalty in the beam search. This increase in penalty continues until a unique ID is produced, or until the diversity penalty reaches a pre-set maximum threshold (e.g., $1$ in this paper). In cases where the maximum penalty is insufficient to generate a unique ID, the algorithm extends the permissible ID length and repeats this process with the initial diversity penalty. Algorithm 1 elaborates on this process.
@@
2:Initialize diversity penalty $\lambda$ to $0.1$
3:Initialize ID length limit $L$ to 1
@@
16:             $\lambda\leftarrow\lambda+0.1$
@@
18:                 Increase $L$ and reset $\lambda$ to $0.1$
@@ 2.4. Base Recommender
After generating item IDs and incorporating them into the prompt template, the system tokenizes the completed prompt and feeds it into the LLM-based recommender. The recommender then generates the tokens of the target recommended item ID in an autoregressive manner. To ensure that the decoded ID corresponds to an actual existing item, we adopt a constrained sequence decoding strategy (De Cao et al., 2020). More specifically, a prefix tree is used to store all generated candidate IDs. Each newly generated token is constrained by the previously generated tokens, ensuring that the generation process only considers tokens that can potentially form an existing candidate ID in the dataset. Suppose the completed input prompt is denoted as a sequence of tokens $\bm{x}=[x_{1},x_{2},\cdots,x_{n}]$. In this case, the base recommender model aims to generate $\bm{y}=[y_{1},y_{2},\cdots,y_{n}]$, where $\bm{y}$ is an ID for a real item in the dataset. We define $\mathcal{V}(y_{<i})$ as the subset of valid tokens in the vocabulary, constrained by the prefix tree with previously generated tokens as nodes. The generation of each token by the decoder is defined as:
@@ 3.1. Datasets
We have $19$ Amazon datasets left, and all of them will be selected to create the massive recommendation corpus for training the foundational model. Since their sizes vary extremely (e.g., Books contains 603,668 users and Automotive only has 2,928 users), we randomly downsample the large datasets to only include $30,000$ users, which is around the median number of users of the Amazon datasets. All the selected recommendation records together form a “Fusion” dataset for training the foundation recommendation model. Complete and detailed data statistics can be seen in Table 3.
@@ 3.2. Implementation Details
In the diverse ID generation algorithm, we set $k=5$ as the number of groups for DBS and start with $\lambda=0.1$ as the initial diversity penalty. This penalty is increased by $0.1$ each iteration when the algorithm fails to generate a unique item ID. If $\lambda$ reaches $1.0$ without producing a valid ID, we then extend the length limit for the item ID. This iterative process of adjusting the diversity penalty and, if necessary, the length of the ID, ensures that the Diverse ID Generation algorithm can successfully produce IDs that are both succinct and distinct. Initially, the token length is set within the range of $[1,10)$, but if the algorithm is unable to generate a unique ID once the diversity penalty has reached its maximum, we increase the token length limit to the range of $[10,20)$. In our experiments, even for item IDs in the largest ID set, specifically the Fusion dataset, only $11.20\%$ of the items required a second attempt at ID generation with an increased diversity penalty, and merely $3.72\%$ of the items necessitated an extension in ID length.
@@ 3.5.1. Baselines
UniSRec (Hou et al., 2022) is the main comparable baseline in the zero-shot setting. It uses an encoder-only structure to learn item representations from their meta text information. We use the inductive version of UniSRec, which has cross-platform capability. In our experiments, we use the same training data for training UniSRec. Although UniSRec has a certain degree of flexibility to learn general information in recommendations, it incorporates some human-inductive knowledge in data preprocessing, e.g., the “title,” “category,” and “brand” information are carefully selected for Amazon datasets. For the Amazon dataset, we still use the pre-selected categories as in the original paper to showcase its best performance, therefore introducing a slight advantage for UniSRec. For Yelp, we use all the plain text information. Since the patterns have not been seen in both models, this is a fair comparison.
@@ 4. Related Works
In recent years, there has been a surge in research on recommendation systems leveraging LLMs. LLM-based recommender systems can be broadly classified into two categories: discriminative and generative (Wu et al., 2023).
@@
In LLM-based discriminative recommendation, LLMs are primarily used to learn better representations of users and items by leveraging contextual information in the recommendation process (Wu et al., 2021; Qiu et al., 2021; Zhang et al., 2022; Yao et al., 2022; Muhamed et al., 2021; Xiao et al., 2022; Hou et al., 2022; Li et al., 2023a). Among them, BERT (Devlin et al., 2018) is commonly used as the backbone. Compared to recommender systems that learn user/item embeddings primarily from user-item associations, these models capture rich textual information in conjunction with collaborative filtering. The core idea of these models is the learning of embeddings with LLMs and integrating the embeddings into a ranking score calculation based paradigm. Since this types of works are not the main focus of this paper, readers may refer to (Wu et al., 2023) for a more comprehensive study.
@@
Our work belongs to the second category: LLM-based generative recommendation (Geng et al., 2022, 2023; Hua et al., 2023; Cui et al., 2022). This novel approach transitions from the traditional ranking-based recommendation to a pure text-to-text method. In comparison with the original score computation and re-ranking framework, this method can directly generate the target item from the complete pool of items (Li et al., 2023b), thus gathering significant attention. Geng et al. (Geng et al., 2022) introduced one of the first text-to-text generative recommendations, leveraging a pre-trained T5 as the recommendation backbone. The model was fine-tuned on $5$ diverse recommendation tasks with predefined prompts and demonstrated promising results on the downstream tasks. Following (Geng et al., 2022), (Geng et al., 2023) proposed a multimodal version with a similar architecture, considering item visual features during recommendation generation. In these works, numerical IDs were assigned to each item, enabling the recommended items to be generated token-by-token. Subsequently, (Hua et al., 2023) posited that indexing item IDs is the most pivotal aspect of generative recommendation. They conducted an extensive study on indexing methods such as sequential indexing, semantic indexing, and collaborative indexing, and assessed their effects on sequential recommendation performance. Although these methods have shown potential, a limitation exists: A primary advantage of using pre-trained language models lies in their capability to understand human language, making it crucial to fully utilize contextual information. However, the rich textual information in recommender platforms is not fully harnessed by using numerical IDs. Furthermore, the employment of Out-Of-Vocabulary (OOV) tokens for item representation limits the generalizability of these methods, confining the trained model to a single dataset. Alternative approaches, such as using item titles (Ji et al., 2024) or categories (Hua et al., 2023) as semantic IDs, may initially seem to offer a more meaningful representation. However, these approaches may encounter duplicated ID issues as the number of items increases, and the IDs may have unintended overlaps. For example, similar titles may represent completely different items. Our method distinguishes itself by proposing an ID-generator that derives semantic item IDs from rich contextual information, enabling both the utilization of this information and facilitating cross-platform, zero-shot recommendations.
@@
Another branch of generative recommendation research has concentrated on probing the potential of LLMs to directly generate recommendations without the need for training or fine-tuning (Liu et al., 2023b; Gao et al., 2023; Liu et al., 2023a; Hou et al., 2023; Sun et al., 2023; Dai et al., 2023). The primary focus in these works is the design of prompts. The emergence of this line of research was inspired by the exceptional zero-shot capabilities of ChatGPT. Researchers are keen to ascertain how ChatGPT performs in a recommendation scenario. However, as pointed out by (Liu et al., 2023b), ChatGPT cannot generate recommendations with accuracy competitive to that of traditional recommendation methods.

### v2

# IDGenRec: LLM-RecSys Alignment with Textual ID Learning

@@ Abstract.
Generative recommendation based on LLMs have transformed the traditional ranking-based recommendation style into a text-to-text generation paradigm. This approach has attracted significant attention. However, in contrast to standard NLP tasks that inherently operate on human vocabulary, current research in generative recommendations struggles to effectively encode recommendation items within the text-to-text framework using concise yet meaningful ID representations. Due to this unresolved issue, the potential of LLM-based generative recommendation systems remains largely unexplored.
@@
Recommender System, Natural Language Processing
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657821isbn: 979-8-4007-0431-4/24/07ccs: Information systems Recommender systemsccs: Computing methodologies Natural language generation
@@ 1. Introduction
A few attempts have been made to tackle this problem. P5 (Geng et al., 2022), one of the first works in generative recommendation, proposes allocating out-of-vocabulary (OOV) tokens to items within the recommendation platform. These assigned IDs are fixed-length numerical tokens roughly created based on their sequential appearance in the dataset (e.g., 1001 for the first item, 1002 for the second item). Later, (Hua et al., 2023) evaluates how recommendation performance can be further improved by using different strategies to initialize the numerical IDs.
@@
We note that some recent LLM-based recommendation models employ an encoder-only structure, similar to BERT, as their representation function. Some of these models (Hou et al., 2022; Li et al., 2023b) also possess (or partially possess) characteristics of a foundational recommendation model, though they do not function as generative models. The distinctions among these models are depicted in Table 1. However, generative models present several advantages over discriminative methods. These include transforming the retrieval and ranking processes into a more streamlined generative process, eliminating the need of one-by-one item score calculation, and leveraging the extensive knowledge embedded in pre-trained generative LLMs. Nonetheless, these encoder-only methods remain valuable as baselines for zero-shot evaluation.
@@ 2.3. Diverse ID Generation
The ID generator employs DBS in generating IDs. To ensure the uniqueness of the generated IDs, the algorithm generates $k$ groups of IDs each time and compares each generated ID against a set of already existing IDs. If a duplicate is detected, the algorithm increases the diversity penalty in the beam search. This increase in penalty continues until a unique ID is produced, or until the diversity penalty reaches a pre-set maximum threshold (e.g., $10$ in this paper). In cases where the maximum penalty is insufficient to generate a unique ID, the algorithm extends the permissible ID length and repeats this process with the initial diversity penalty. Algorithm 1 elaborates on this process.
@@
2:  Initialize diversity penalty $\lambda$ to $1$
3:  Initialize ID length limit $L$ to 10
@@
16:           $\lambda\leftarrow\lambda+1$
@@
18:              Increase $L$ and reset $\lambda$ to $1$
@@ 2.4. Base Recommender
After generating item IDs and incorporating them into the prompt template, the system tokenizes the completed prompt and feeds it into the LLM-based recommender. The recommender then generates the tokens of the target recommended item ID in an autoregressive manner.
To ensure that the decoded ID corresponds to an actual existing item, we adopt a constrained sequence decoding strategy (De Cao et al., 2020). More specifically, a prefix tree is used to store all generated candidate IDs. Each newly generated token is constrained by the previously generated tokens, ensuring that the generation process only considers tokens that can potentially form an existing candidate ID in the dataset. Suppose the completed input prompt is denoted as a sequence of tokens $\bm{x}=[x_{1},x_{2},\cdots,x_{n}]$. In this case, the base recommender model aims to generate $\bm{y}=[y_{1},y_{2},\cdots,y_{n}]$, where $\bm{y}$ is an ID for a real item in the dataset. We define $\mathcal{V}(y_{<i})$ as the subset of valid tokens in the vocabulary, constrained by the prefix tree with previously generated tokens as nodes. The generation of each token by the decoder is defined as:
@@ 3.1. Datasets
We have $19$ Amazon datasets left, which are used to create the massive recommendation corpus for training the foundational model. Since their sizes vary extremely (e.g., Books contains 603,668 users and Automotive only has 2,928 users), we randomly downsample the large datasets to only include $30,000$ users, which is around the median number of users of the Amazon datasets. All the selected recommendation records together form a “Fusion” dataset for training the foundation recommendation model. Complete and detailed data statistics can be seen in Table 3.
@@ 3.2. Implementation Details
In the diverse ID generation algorithm, we set $k=10$ as the number of groups for DBS and start with $\lambda=1$ as the initial diversity penalty. This penalty is increased by $1$ in each iteration until the algorithm successfully generates a unique item ID. If $\lambda$ reaches $10$ without producing a valid ID, we then extend the length limit for the item ID. This iterative process of adjusting the diversity penalty and, if necessary, the length of the ID, ensures that the Diverse ID Generation algorithm can successfully produce IDs that are both succinct and distinct. Initially, the token length is set within the range of $[1,10)$, but if the algorithm is unable to generate a unique ID once the diversity penalty has reached its maximum, we increase the token length limit to the range of $[10,20)$. In our experiments, even for item IDs in the largest ID set, specifically the Fusion dataset, only $11.20\%$ of the items required a second attempt at ID generation with an increased diversity penalty, and merely $3.72\%$ of the items necessitated an extension in ID length.
@@ 3.5.1. Baselines
UniSRec (Hou et al., 2022) is the main comparable baseline in the zero-shot setting, which uses an encoder-only structure to learn item representations from their meta text information. We note that UniSRec incorporates some human-inductive knowledge in data preprocessing, e.g., the “title,” “category,” and “brand” information are carefully selected for Amazon datasets. For the Amazon dataset, we still use the pre-selected categories as in the original paper to showcase its best performance, therefore introducing a slight advantage for UniSRec. For Yelp, we use all the plain text information. Since the patterns have not been seen in both models, this is a fair comparison.
@@ 4. Related Works
Recent LLM-based recommender systems can be broadly classified into two categories: discriminative and generative (Wu et al., 2023).
@@
In LLM-based discriminative recommendation, LLMs are primarily used to learn better representations of users and items by leveraging contextual information in the recommendation process (Wu et al., 2021; Qiu et al., 2021; Zhang et al., 2022; Yao et al., 2022; Muhamed et al., 2021; Xiao et al., 2022; Hou et al., 2022; Li et al., 2023b; Yuan et al., 2023; Li et al., 2023a). Among them, BERT (Devlin et al., 2018) is commonly used as the backbone. Compared to recommender systems that learn user/item embeddings primarily from user-item associations, these models capture rich textual information in conjunction with collaborative filtering. The core idea of these models is the learning of embeddings with LLMs and integrating the embeddings into a ranking score calculation based paradigm. Since this types of works are not the main focus of this paper, readers may refer to (Wu et al., 2023) for a more comprehensive study.
@@
Our work belongs to the second category: LLM-based generative recommendation (Geng et al., 2022, 2023; Hua et al., 2023; Cui et al., 2022). This novel approach transitions from the traditional ranking-based recommendation to a pure text-to-text method. In comparison with the original score computation and re-ranking framework, this method can directly generate the target item from the complete pool of items (Li et al., 2023c), thus gathering significant attention. Geng et al. (Geng et al., 2022) introduced one of the first text-to-text generative recommendations, leveraging a pre-trained T5 as the recommendation backbone. The model was fine-tuned on $5$ diverse recommendation tasks with predefined prompts and demonstrated promising results on the downstream tasks. Following (Geng et al., 2022), (Geng et al., 2023) proposed a multimodal version with a similar architecture, considering item visual features during recommendation generation. In these works, numerical IDs were assigned to each item, enabling the recommended items to be generated token-by-token. Subsequently, (Hua et al., 2023) posited that indexing item IDs is the most pivotal aspect of generative recommendation. They conducted an extensive study on indexing methods and assessed their effects on sequential recommendation performance. Although these methods have shown potential, a limitation exists: The rich textual information in recommender platforms is not fully harnessed by using numerical IDs. Furthermore, the employment of OOV tokens for item representation limits the generalizability of these methods, confining the trained model to a single dataset. Alternative approaches, such as using item titles (Ji et al., 2024) or categories (Hua et al., 2023) as semantic IDs, may initially seem to offer a more meaningful representation. However, these approaches may encounter duplicated ID issues as the number of items increases, and the IDs may have unintended overlaps. For example, similar titles may represent completely different items. Our method distinguishes itself by proposing an ID-generator that derives semantic item IDs from rich contextual information, enabling both the utilization of this information and facilitating cross-platform, zero-shot recommendations.
@@
Another branch of generative recommendation research has concentrated on probing the potential of LLMs to directly generate recommendations without the need for training or fine-tuning (Liu et al., 2023b; Gao et al., 2023; Liu et al., 2023a; Hou et al., 2023; Sun et al., 2023; Dai et al., 2023). The primary focus in these works is the design of prompts. The emergence of this line of research was inspired by the exceptional zero-shot capabilities of ChatGPT, and they aim to discover how ChatGPT performs in a recommendation scenario. However, as pointed out by (Liu et al., 2023b), ChatGPT cannot generate recommendations with accuracy competitive to that of traditional recommendation methods.